# The Unbearable Slowness of Being

### v1

@@ 6.3 Complexity bottleneck
Psychological treatments of this subject have invoked a number of metaphors for the serial nature of human cognition, such as “single channel operation” [welford_single-channel_1967], “attentional bottleneck” [broadbent_perception_1958], and “limited processing resources” [norman_data-limited_1975]. These narratives share a common framework: Sensory systems collect high-dimensional signals at a high rate with many pathways operating in parallel. After considerable filtering and reduction, these signals must compete for some central neural resource, where goals, emotions, memory, and sensory data are combined into low-dimensional decisions. This central neural resource can only execute its functions in series, which constrains the low throughput. But these are not explanations unless one specifies the identity of that limiting neural resource, and to date there aren’t any convincing proposals.
@@
The types of cognitive functions that have been associated with the attentional bottleneck do not seem to place high demands on neural hardware. We know this from explicit simulations of neural circuits. For example, wang_decision_2008 built a realistic circuit model of decision-making that replicates many of the phenomena observed in perceptual psychophysics experiments on macaques. The circuit contains just 2000 integrate-and-fire neurons, and there is no reason to think that this is a lower limit. Hundreds of these decision-making modules would fit into just one square millimeter of cortex.
@@
A much more demanding cognitive task is image recognition. A neural network model called AlexNet was trained to classify about 1 million photographs into 1000 object categories [krizhevsky_imagenet_2012, rawat_deep_2017]. That means extracting 10 bits from an image with 1.2 million bits. This network used 650,000 neurons, the equivalent of just a few square millimeters of cortex. Again, it was not designed to minimize the number of units.  [^2] For an implementation with integrate-and-fire units, see [hunsberger_spiking_2015]. It is difficult to see why we should not have many of these networks running in parallel, rather than a single central pathway for image recognition.
@@
Another useful point of reference: A fruit fly has less than 200,000 neurons [raji_number_2021]. Somehow that is sufficient for all the complex operations, including visually-controlled aerobatics, olfactory navigation while walking or flying, social communication, mating and aggression, and more. The human prefrontal cortex alone contains enough neuronal hardware to run 5,000 flies. Why can it not implement a few tasks to run simultaneously?
The mystery of “limiting neural resources” is particularly vexing because the cerebral cortex is touted as the substrate for both rapid and parallel sensory processing (e.g. visual cortex) and for slow and serial cognition (e.g. the prefrontal cortex). However, different areas of the human cortex have essentially the same cell types and circuit architecture. Are we to believe that the billion neurons in the primary visual cortex elaborate many channels of visual processing in 10,000 parallel modules, whereas the billion neurons in the prefrontal cortex just deal with a handful of slow variables, like rules and values, associated with the single task currently at hand [miller_prefrontal_2000]?
@@
It seems that we need to elaborate our model of what these more cognitive brain areas are doing. The current understanding is not commensurate with the enormous processing resources available, and we have seen no viable proposal for what would create a neural bottleneck that forces single-strand operation. We will return to this dilemma in the last section.
6.3.1 Inefficient neural hardware
Another argument one hears frequently is that neurons are simply inefficient devices for information processing, so nature needs to employ huge numbers of them to complete even simple operations. In particular, individual neurons – unlike transistors – are subject to random biochemical noise, so perhaps one needs to average over many neurons to get a reliable signal. Also, perhaps the brain includes a large amount of redundancy, with populations of essentially identical neurons, to guard against the loss of neurons from aging or injury. Neither of these explanations seems plausible.
Regarding “noisy neurons”, the information-theoretic analysis reviewed above (section 3) gets at exactly that point and shows that a single neuron can reliably signal several bits per spike about its dendritic inputs. As a rule, once experimenters discover what a particular neuron encodes, they realize that the representation is amazingly precise. A single optic nerve fiber can reliably signal the arrival of its preferred feature in the visual stimulus with a precision of 1 millisecond [berry_structure_1997]. Even deep in the primate cortex, a single neuron can distinguish between two visual stimuli as precisely as the entire monkey [britten_analysis_1992]. There is increasing recognition in the community that what we used to call noise in neural recordings is actually a reliable signal about variables that the experimenter doesn’t control [ringach_spontaneous_2009, stringer_spontaneous_2019].
On the subject of redundancy, again there is precious little evidence for duplication of essentially identical neurons. For example, in the fruit fly, many neuron types exist in just two copies, one on each side of the brain [sterne_classification_2021, aso_mushroom_2014]. One might argue that fruit flies are cheap and short-lived, so nature simply duplicates the entire organism to ensure robustness. However, the principle applies even to large animals like ourselves. In the primate retina, every point in the visual field is covered by just one neuron of each cell type, with very little overlap [sanes_types_2015, wassle_cone_2009, reese_design_2015]. In the visual cortex, a small stroke can abolish vision in a section of the visual field. Apparently, there is no redundancy built in for one of the most common forms of brain injury.
@@ 7 Outer brain vs inner brain
Here we consider some promising research directions related to the central paradox. To recapitulate, the discrepancy between peripheral processing and central cognition suggests that the brain operates in two distinct modes: The “outer brain” is closely connected to the external world through sensory inputs and motor outputs. This is a realm of high dimensionality: many millions of sensory receptors and muscle fibers, and the information rates are extremely high. The “inner brain”, on the other hand, operates on a dramatically reduced data stream, filtered to the essential few bits that matter for behavior at any one moment. The challenge for the inner brain is to combine the animal’s goals with current inputs from the world and previous memories to make decisions and trigger new actions. The information rates are very low, but the processing must remain flexible because context and goals can shift at a moment’s notice. A number of interesting research questions emerge regarding the relationship between the inner and outer brain.  [^3] Note that inner/outer are not anatomical terms, but reflect the relation to the sensory/motor periphery.
@@
First, how can the inner and the outer brain communicate? The slow inner brain listening to the sensory onslaught from the outer brain seems like ‘drinking from the Hoover Dam’.  [^4] The rate of water flow through the Hoover Dam is $\sim 10^{8}$ times the rate of human drinking, the same ratio as for information rates in the outer vs the inner brain. Presumably, the matching occurs in multiple steps, along some graded continuum of information rate. For example, at the output of the retina, the image information has already been reduced by a factor of 10 or more, leaving only image features that are useful in downstream processing [roska_retina_2014, field_information_2007, wassle_parallel_2004]. The retina projects directly to the superior colliculus (SC, also called optic tectum), which is thought to act as a further bottleneck. Neurons in the upper layers of the SC have fine-grained receptive fields similar to those in the retina. By contrast, the lower layers of the SC represent pre-motor signals that reflect the animal’s behavioral decisions. Examples include the saccadic eye movements of monkeys, reorienting of the head in owls, or approach vs avoidance of a novel object by rodents  [basso_unraveling_2021, knudsen_instructed_2002, may_mammalian_2006]. Thus it appears that the SC distills a massively parallel and high-dimensional sensory input into a map of suitable targets for action. Importantly, the circuit mechanisms underlying this “sifting” of visual information are coming into view [branco_neural_2020, lee_sifting_2020, de_malmazet_collicular_2023]. Recent work suggests that the SC also plays a role in phenomena of attention that do not involve overt movement [krauzlis_superior_2013]. Perhaps it contributes to a bottom-up “saliency map” [itti_saliency-based_2000] that ultimately directs the attentional bottleneck in cognition.
@@
Second, what are the principles of neural function on the two sides of the interface? Taking the water flow analogy further: The design of the Hoover Dam turbines relies on rather different engineering principles from the design of a beer bottle. No one would consider making them of the same materials, given the enormous difference in performance. Yet the brain seems to rely on the same materials throughout: neurons, synapses, and glia. Very similar neocortical circuits seem to be engaged in both modes. Are we missing some principles of brain design that categorically differentiate the functions of the inner and outer brain?
@@
Comparing research reports from the two sides of the inner/outer brain divide can be a challenge, because the practitioners tend to operate on entirely different assumptions. In the sensory regions, tradition holds that every spike is sacred, because each neuron has its own receptive field, and single spikes are known to convey a few bits of information each about the stimulus [hubel_early_1998, callaway_local_1998, fitzpatrick_seeing_2000]. For example, the retina has one million output fibers, each with its own visual receptive field. These signals are transmitted with little loss to the primary visual cortex, so one can reasonably argue that the dimensionality of neural activity in V1 is 1 million. By contrast, researchers working closer to behavioral output, say in the prefrontal cortex or motor cortex, are often happy to boil the activity of millions of neurons down to just two or three “latent dimensions” [gallego_neural_2017, durstewitz_reconstructing_2023, jazayeri_interpreting_2021]. The typical finding identifies a low-dimensional manifold on which the neuronal population vector travels, following simple dynamics. We see here that the paradoxical contrast between inner and outer brain information rates translates to an equally stark contrast regarding the presumed dimensionality of neural activity. It seems difficult to accept that two regions of the neocortex with overtly similar architecture (ignoring the minor differences [hilgetag_natural_2022]) are organized to operate in such radically different ways.
@@
Alternatively, it is conceivable that the empirically observed difference in dimensionality is a consequence of experimental design. Most of the studies of inner brain phenomena involve experimental tasks of very low complexity, like a mouse making repeated binary choices between the same two stimuli [burgess_high-yield_2017], or a monkey moving a lever along a few possible directions [shenoy_cortical_2013]. Obviously, the resulting neural representations will be no more complex than what the animal is doing, especially if one averages over many trials [gao_simplicity_2015]. Under those conditions, one finds by necessity that some low-dimensional manifold accounts for much of neuronal dynamics. However, real-life behavior is not that simple. For example, the Speed Card player, shuffling through the card deck, switches between different neural computations every few tenths of a second: Every eye saccade brings a new card image in focus and triggers a bout of image recognition. These episodes are interleaved with storage in short-term memory, and again with an update of the inner narrative or memory palace that can hold the new card, followed by another saccade, all within a second or less. In conversations, we switch quickly between listening and talking, sometimes interleaved with thinking for a fraction of a second. When driving, we check the windshield, dashboard, rear mirror, and side mirrors every few seconds, and process the results in entirely different modes, like estimating distance from the road edge vs reading highway signs. Thus, we really perform thousands of different “microtasks” on any given day, switching between them as fast as we can saccade, up to a few times per second [land_roles_1999, hayhoe_visual_2003].
@@
On this background, it would be informative to study the inner brain under such naturalistic conditions that involve a rapid sequence of distinct microtasks [gallego_neural_2017]. Without experimental designs tuned to the specialty of the inner brain, we might fail altogether to discover its essential mechanism. Imagine a counterfactual history in which our understanding of the primary visual cortex (V1) was purely based on low-dimensional stimuli: We trained a macaque monkey to fixate while viewing a large rotating propeller that reversed direction periodically. After recording from thousands of neurons in the monkey’s V1, we discovered that most of the cells responded to this stimulus, so we took a dimension-reduction approach. Fortunately, much of the population activity was captured with just two principal components. And in those reduced dimensions the population vector had rotational dynamics. Finally, we decoded the direction of rotation of the propeller from the rotation of the neural trajectory. After three years of peer review the paper appeared in a high-impact journal.
@@
Given what we know today, none of these hypothetical results touch even remotely on what the visual cortex does. It was only through fine-grained high-dimensional stimuli and single-neuron analysis that researchers uncovered the high-resolution structure of receptive fields, the functional anatomy of their arrangement, and the circuit mechanisms in V1  [hubel_early_1998, callaway_local_1998, carandini_we_2005, gegenfurtner_cortical_2003, ohki_specificity_2007]. Returning to the prefrontal cortex – which contains about the same number of neurons as V1 [gabiNoRelativeExpansion2016, dorph-petersenPrimaryVisualCortex2007] – it is conceivable that we are missing a similar fine-grained organization. Could there be thousands of small modules in the prefrontal cortex, each dedicated to a specific microtask, much as a hypercolumn in V1 specializes for a small patch of the visual field [mountcastle_columnar_1997]? This is very speculative, but the only way to resolve our ignorance here is to perform new kinds of experiments.
@@
Appendix A Measuring the information rate of human behavior
@@
  • Calculate the entropy according to Eqn 2.
However, that result is sensitive to how finely I measure your actions. As you scratch your toe, should I measure that movement in microns or centimeters? The finer I sample the movements to distinguish different actions $a_{i}$, the larger the entropy according to Eqn 2. So what is a natural bin size for actions?
@@
So the presence of a task leads to a natural binning of the actions, in this case into keystrokes. It also allows us to separate the signal, namely the entropy that determines task performance, from noise, namely the part that does not. The difference between total entropy and noise entropy is the mutual information
@@
Appendix B The information rate of human behavior: more examples
@@
B.1 Video games and motor capacity
@@
The early video games are good case studies, as it is possible to enumerate all possible outcomes based on the discretized graphics and limited types of button presses. In the game Tetris, for example, players are required to stack different shapes of four-unit pieces (called tetrominoes) being dropped at a constant rate, one at a time from the top of the screen. To prevent the pieces from accumulating, players have to arrange the pieces to fill each ten-unit row that will be removed once completed. To estimate the upper bound of information, we can assume that every placement of a tetromino – its orientation and location in the row – is equally probable. Of all the tetrominoes, the T-shape piece has the highest number of possible orientations (4) $\times$ locations (9) to choose from. Multiply this by the highest human placement rate (1.27 pieces per second on average for the highest tier gamers [tenchi_tetrio_2021]), and you will get an information rate lower than 6 bits/s.
@@
For convenient comparisons across games, we can measure the prowess of a player by actions per minute (APM) to estimate information rates. While the best Tetris players perform at 26.63 APM on average [tenchi_tetrio_2021], professional StarCraft players reach 1000 APM during intense battle sequences [guinness_world_records_limited_most_2023]. Again there is a considerable amount of redundancy compared to typing English, because only a few important hotkeys are frequently used for the game, and many presses serve the same given order. Thus, the entropy of the keystroke string in the game is very likely to be lower than that of English letter strings (1 bit/character). We will generously credit the StarCraft champion with zero errors and 1 bit/character as the typist, and conclude that the information throughput at the very height of battle is at the same order of magnitude as a typist during a regular work day, at $\sim$ 16.7 bits/s.
@@
Granted, our fingers did not evolve for typing, and one may argue that the human motor system can be measured by tasks with more freedom of movement than button bashing. fitts_information_1954 trained the participants to perform four different types of motor tasks that involved relocating one’s forearms rapidly between confined target areas. At their best performance within a short spree of 20 seconds, the well-practiced subjects reached 10-12 bits/s in all four tasks. Different speed or accuracy requirements yielded an invariant human motor capacity at 10 bits/s [fitts_information_1964].
B.2 Language

### v2

@@ 7.4 Complexity bottleneck
Psychological treatments of this subject have invoked a number of metaphors for the serial nature of human cognition, such as “single channel operation” welford_single-channel_1967, “attentional bottleneck” broadbent_perception_1958, zhaoping_peripheral_2023, and “limited processing resources” norman_data-limited_1975. These narratives share a common framework: Sensory systems collect high-dimensional signals at a high rate with many pathways operating in parallel. After considerable filtering and reduction, these signals must compete for some central neural resource, where goals, emotions, memory, and sensory data are combined into low-dimensional decisions. This central neural resource can only execute its functions in series, which constrains the low throughput. But these are not explanations unless one specifies the identity of that limiting neural resource, and to date there aren’t any convincing proposals.
@@
The types of cognitive functions that have been associated with the attentional bottleneck do not seem to place high demands on neural hardware. We know this from explicit simulations of neural circuits. For example, Wang wang_decision_2008 built a realistic circuit model of decision-making that replicates many of the phenomena observed in perceptual psychophysics experiments on macaques. The circuit contains just 2000 integrate-and-fire neurons, and there is no reason to think that this is a lower limit. Hundreds of these decision-making modules would fit into just one square millimeter of cortex.
@@
A much more demanding cognitive task is image recognition. A neural network model called AlexNet was trained to classify about 1 million photographs into 1000 object categories krizhevsky_imagenet_2012, hunsberger_spiking_2015, rawat_deep_2017. That means extracting 10 bits from an image with 1.2 million bits. This network used 650,000 neurons, the equivalent of just a few square millimeters of cortex. Again, it was not designed to minimize the number of units. It is difficult to see why we should not have many of these networks running in parallel, rather than a single central pathway for image recognition.
@@
Another useful point of reference: A fruit fly has less than 200,000 neurons raji_number_2021. Somehow that is sufficient for all the complex operations, including visually-controlled aerobatics, olfactory navigation while walking or flying, social communication, mating and aggression, and more. The human prefrontal cortex alone contains enough neuronal hardware to run 5,000 flies. Why can it not implement a few tasks to run simultaneously? What are the ‘limiting neural resources”?
@@
It seems that we need to elaborate our model of what these more cognitive brain areas are doing. The current understanding is not commensurate with the enormous processing resources available, and we have seen no viable proposal for what would create a neural bottleneck that forces single-strand operation. We elaborate on this dilemma in the last section.
@@ 8 Outer brain vs inner brain
The discrepancy between peripheral processing and central cognition suggests that the brain operates in two distinct modes: The “outer brain” is closely connected to the external world through sensory inputs and motor outputs. This is a realm of high dimensionality: many millions of sensory receptors and muscle fibers, and extremely high information rates. The “inner brain”, on the other hand, operates on a dramatically reduced data stream, filtered to the essential few bits that matter for behavior at any one moment. The challenge for the inner brain is to combine the animal’s goals with current inputs from the world and previous memories to make decisions and trigger new actions. The information rates are very low, but the processing must remain flexible because context and goals can shift at a moment’s notice. A number of interesting research questions emerge regarding the relationship between the inner and outer brain.
First, how can the inner and the outer brain communicate? The slow inner brain listening to the sensory onslaught from the outer brain seems like ‘drinking from the Hoover Dam’. The rate of water flow through the Hoover Dam is $\sim 10^{8}$ times the rate of human drinking, the same ratio as between information rates in the outer vs the inner brain (Eqn 2). Presumably, the matching occurs in multiple steps, along some graded continuum of information rate. For example, at the output of the retina, the image information has already been reduced by a factor of 10 or more, leaving only image features that are useful in downstream processing roska_retina_2014, field_information_2007, wassle_parallel_2004. Given the large sifting number (Eqn 2), there are many log units of visual compression that remain to be understood.
@@
The retina projects directly to the superior colliculus (SC, also called optic tectum), which is thought to act as a further bottleneck. Neurons in the upper layers of the SC have fine-grained receptive fields similar to those in the retina. By contrast, the lower layers of the SC represent pre-motor signals that reflect the animal’s behavioral decisions. Examples include the saccadic eye movements of monkeys, reorienting of the head in owls, or approach vs avoidance of a novel object by rodents  basso_unraveling_2021, knudsen_instructed_2002, may_mammalian_2006. Thus it appears that the SC distills a massively parallel and high-dimensional sensory input into a map of suitable targets for action. The circuit mechanisms underlying this sifting are coming into view branco_neural_2020, lee_sifting_2020, de_malmazet_collicular_2023. Recent work suggests that the SC also plays a role in phenomena of attention that do not involve overt movement krauzlis_superior_2013. Perhaps it contributes to a bottom-up “saliency map” itti_saliency-based_2000 that ultimately directs the attentional bottleneck in cognition.
@@
Second, what are the principles of neural function on the two sides of the interface? Taking the water flow analogy further: The design of the Hoover Dam turbines relies on rather different engineering principles from the design of a beer bottle. No one would consider making them of the same materials, given the enormous difference in performance. Yet the brain seems to rely on the same materials throughout: neurons, synapses, and glia. The cerebral cortex is touted as the substrate for both rapid and parallel sensory processing (e.g. visual cortex) and for slow and serial cognition (e.g. the prefrontal cortex). Very similar neocortical circuits seem to be engaged in both modes. Are we missing some principles of brain design that categorically differentiate the functions of the inner and outer brain?
@@
Comparing research reports from the two sides of the inner/outer brain divide can be a challenge, because the practitioners tend to operate on entirely different assumptions. In the sensory regions, tradition holds that “every spike is sacred”, because each neuron has its own receptive field, and single spikes are known to convey a few bits of information each about the stimulus hubel_early_1998, callaway_local_1998, fitzpatrick_seeing_2000. For example, the retina has one million output fibers, each with its own visual receptive field. These signals are transmitted with little loss to the primary visual cortex, so one can reasonably argue that the dimensionality of neural activity in V1 is 1 million.
By contrast, researchers working closer to behavioral output, say in the prefrontal cortex or motor cortex, are often happy to boil the activity of millions of neurons down to just two or three “latent dimensions” gallego_neural_2017, durstewitz_reconstructing_2023, jazayeri_interpreting_2021. The typical finding identifies a low-dimensional manifold on which the neuronal population vector travels, following simple dynamics. We see here that the paradoxical contrast between inner and outer brain information rates translates to an equally stark contrast regarding the presumed dimensionality of neural activity. Are we to believe that the billion neurons in the primary visual cortex elaborate many channels of visual processing in 10,000 parallel modules, while the billion neurons in the prefrontal cortex just deal with a handful of slow variables, like rules and values, associated with the single task currently at hand miller_prefrontal_2000. It seems difficult to accept that two regions of the neocortex with overtly similar cell types and architecture (ignoring the minor differences hilgetag_natural_2022) are organized to operate in such radically different ways.
@@
Alternatively, the empirically observed difference in dimensionality may be an artifact of experimental design. Most of the studies of inner brain phenomena involve experimental tasks of very low complexity, like a mouse making repeated binary choices between the same two stimuli burgess_high-yield_2017, or a monkey moving a lever along a few possible directions shenoy_cortical_2013. Obviously, the resulting neural representations will be no more complex than what the animal is doing, especially if one averages over many trials gao_simplicity_2015. Under those conditions, one finds by necessity that some low-dimensional manifold accounts for much of neuronal dynamics. However, real-life behavior is not that simple. For example, the Speed Card player, shuffling through the card deck, switches between different neural computations every few tenths of a second: Every eye saccade brings a new card image in focus and triggers a bout of image recognition. These episodes are interleaved with storage in short-term memory, and again with an update of the inner narrative or memory palace that can hold the new card, followed by another saccade. In conversations, we switch quickly between listening and talking, sometimes interleaved with a moment of thinking. When driving, we check the windshield, dashboard, rear mirror, and side mirrors, and process the results in entirely different modes, like estimating distance from the road edge vs reading highway signs. The act of making tea requires 45 different brief subtasks land_eye_2006. Thus, we really perform thousands of different “microtasks” on any given day, switching between them as fast as we can saccade land_roles_1999, hayhoe_visual_2003. Each of the microtasks depends on real-time feedback, requiring the sifting of sensory inputs and expansion into motor outputs within sub-second timeframes. The flexible configuration and control of all these data streams seems essential to our cognitive functions. Perhaps the associated routing machinery accounts for the billions of neurons in the inner brain?
@@
On this background, it would be informative to study the inner brain under such naturalistic conditions where it controls a rapid sequence of distinct microtasks gallego_neural_2017. Without experimental designs tuned to the specialty of the inner brain, we might fail altogether to discover its essential mechanism. Imagine a counterfactual history in which our understanding of the primary visual cortex (V1) was purely based on low-dimensional stimuli: “We trained a macaque monkey to fixate while viewing a large rotating propeller that reversed direction periodically. After recording from thousands of neurons in the monkey’s V1, we discovered that most of the cells responded to this stimulus, so we took a dimension-reduction approach. Fortunately, much of the population activity was captured with just two principal components. And in those reduced dimensions the population vector had rotational dynamics. Finally, we decoded the direction of rotation of the propeller from the rotation of the neural trajectory. After three years of peer review the paper appeared in a high-impact journal.”
@@
Given what we know today, none of these hypothetical results touch even remotely on what the visual cortex does. It was only through fine-grained high-dimensional stimuli and single-neuron analysis that researchers uncovered the high-resolution structure of receptive fields, the functional anatomy of their arrangement, and the circuit mechanisms in V1  hubel_early_1998, callaway_local_1998, carandini_we_2005, gegenfurtner_cortical_2003, ohki_specificity_2007. Returning to the prefrontal cortex – which contains about the same number of neurons as V1 gabiNoRelativeExpansion2016, dorph-petersenPrimaryVisualCortex2007 – it is conceivable that we are missing a similar fine-grained organization. Could there be thousands of small modules in the prefrontal cortex, each dedicated to a specific microtask, much as a hypercolumn in V1 specializes for a small patch of the visual field mountcastle_columnar_1997? This is very speculative, but the only way to resolve our ignorance here is to perform new kinds of experiments.
@@
Box: Measuring the information rate of human behavior
@@
  • Calculate the entropy according to Eqn 3.
However, that result is sensitive to how finely I measure your actions. As you scratch your toe, should I measure that movement in microns or centimeters? The finer I sample the movements to distinguish different actions $a_{i}$, the larger the entropy according to Eqn 3. So what is a natural bin size for actions?
@@
So the presence of a concrete task leads to a natural binning of the actions, in this case into keystrokes. It also allows us to separate the signal, namely the entropy that determines task performance, from noise, namely the part that does not. The difference between total entropy and noise entropy is the mutual information
@@
We adopt the use of information theoretic measures here, because they allow a comparison of behavioral performance in neurons, networks, entire organisms, and machines using the same units of bits per second. In simple terms, a behaving agent chooses one among many possible actions given a set of inputs, and the information rate measures how much that range of actions gets constrained by the decisions. Information theory is not the only statistical framework one could employ, and we don’t consider it a magic bullet shannon_bandwagon_1956.
Appendix A The information rate of human behavior: more examples
@@
A.1 Video games and motor capacity
@@
The early video games are good case studies, as it is possible to enumerate all possible outcomes based on the discretized graphics and limited types of button presses. In the game Tetris, for example, players are required to stack different shapes of four-unit pieces (called tetrominoes) being dropped at a constant rate, one at a time from the top of the screen. To prevent the pieces from accumulating, players have to arrange the pieces to fill each ten-unit row that will be removed once completed. To estimate the upper bound of information, we can assume that every placement of a tetromino – its orientation and location in the row – is equally probable. Of all the tetrominoes, the T-shape piece has the highest number of possible orientations (4) $\times$ locations (9) to choose from. Multiply this by the highest human placement rate (3-4 pieces per second on average for the highest tier gamers tetrio_tetra_2024), and you will get an information rate of around 7 bits/s.
@@
For convenient comparisons across games, we can measure the prowess of a player by actions per minute (APM) to estimate information rates. While the best Tetris players perform at $\sim$200 APM tetrio_tetra_2024, professional StarCraft players reach 1000 APM during intense battle sequences guinnessworldrecordslimited_most_2023. Again there is a considerable amount of redundancy compared to typing English, because only a few important hotkeys are frequently used for the game, and many presses serve the same given order. Thus, the entropy of the keystroke string in the game is very likely to be lower than that of English letter strings (1 bit/character). We will generously credit the StarCraft champion with zero errors and 1 bit/character as the typist, and conclude that the information throughput at the very height of battle is at the same order of magnitude as a typist during a regular work day, at $\sim$ 16.7 bits/s.
@@
Granted, our fingers did not evolve for typing, and one may argue that the human motor system can be measured by tasks with more freedom of movement than button bashing. Fitts fitts_information_1954 trained the participants to perform four different types of motor tasks that involved relocating one’s forearms rapidly between confined target areas. At their best performance within a short spree of 20 seconds, the well-practiced subjects reached 10-12 bits/s in all four tasks. Different speed or accuracy requirements yielded an invariant human motor capacity at 10 bits/s fitts_information_1964.
A.2 Language